Magnetic Field. I have no idea how to solve these questions, please help!!!!
Answer #1

formula for magnetic field of circular coil is

B = uo NI/ 2R

B/I = uo N/2 R

= 4pi * 10^-7 ( 100)/2 ( 0.105 m)

=5.98 * 10^-4 T/A

(b)

if B/I = 6.3 * 10^-4 T/A

B = uo NI/ 2R

uo = B ( 2R)/NI

= 2 ( 0.105)(6.3 * 10^-4 T/A)/100

=1.32 * 10^-6

(3)

e= A dB/ dt

= pi ( 0.105)^2 ( 0.3 * 10^-3 T)/0.05

=0.207 mV or 0.207* 10^-3 V